OpenStudy (anonymous):

John and Alan have a collection of x baseball cards. John has x/4 cards. What fraction of the cards does Alan have?

OpenStudy (kropot72):

Alan has x - x/4 cards. The required fraction is \[\frac{x-\frac{x}{4}}{x}\] Now you need to simplify.

OpenStudy (anonymous):

he has 3/4

OpenStudy (kropot72):

Good work. Your answer is correct.

OpenStudy (anonymous):

do you know Simplify -1 3/7x-3 2/3

OpenStudy (kropot72):

\[-1\frac{3}{7}\times -3\frac{2}{3}\] Is this the question?

OpenStudy (anonymous):

yes but the (-3 2/3) is supposed to be like that sorry

OpenStudy (kropot72):

The first step is to make both the mixed numbers into fractions. Can you do that?

OpenStudy (anonymous):

no its been so long

OpenStudy (anonymous):

im really bad at fractions

OpenStudy (kropot72):

-1 3\7 = -10/7 -3 2/3 = -11/3 \[-\frac{10}{7}\times -\frac{11}{3}=?\]

OpenStudy (anonymous):

2 12/30??

OpenStudy (kropot72):

\[-\frac{10}{7}\times -\frac{11}{3}=\frac{-10\times -11}{7\times 3}=?\]

OpenStudy (anonymous):

21/21

OpenStudy (anonymous):

is thee answer 5 5/21

OpenStudy (kropot72):

Yes, your last answer is correct.

OpenStudy (anonymous):

yay

OpenStudy (kropot72):

Well done :)